A jetliner flying with the wind traveled 3600 km in 6 hours. The return trip, against the wind, took 8 hours. Find the rate at which the jet flew in still air.
Thanks

Answer by checkley77(12844)

3600/6=600 mph with the wind.
3600/8=450 mph against the wind.
(600+450)/2=1050/2=525 is the ave. speed in still air.